A plane which is flying horizontally at a constant speed Vo and at a height h above the sea must drop a bundle of supplies to a castaway on a small raft. Write down newtons second law for the bundle as it falls from the plane, assuming you can neglect air resistance.

1
Expert's answer
2022-02-25T11:14:23-0500

Newton's 2nd law,

F=ma.

F/m=dv/dt

v=(F/m)t + c

At t=0, v=v0, then

c=v0 ds/dt=(F/m)t + v0

s=0.5(F/m)t2 + v0t + c

Taking at t=0, s=0 then c=0 s=0.5(F/m)t2 + v0t